OpenStudy (anonymous):

Can someone help please In 1993, the life expectancy of males in a certain country was 70.8 years. In 1997, it was 73.3 years. Let E represent the life expectancy in year t and let t represent the number of years since 1993. The linear function E(t) that fits the data is E(t) = -- + -- round to nearest tenth Use the function to predict the life expectancy of males in 2007. E(14) = --- round to the nearest tenth

OpenStudy (bahrom7893):

hmm... So I will let years be the x - axis or the variable, and the life expectancy will be y.. x = 3, y = 70.8 x = 7, y = 73.3

OpenStudy (bahrom7893):

Now write the equation using those two points: y-y1 = M(x-x1), where M is the slope.

OpenStudy (bahrom7893):

M = (y-y1)/(x-x1) = (73.3-70.8)/(7-3) = 2.5/4

OpenStudy (bahrom7893):

M = 0.625 so: y-73.3 = 0.625(x-7) y = 0.625x - 0.625*7 + 73.3

OpenStudy (bahrom7893):

y = 0.625x + 68.95

OpenStudy (bahrom7893):

or you could just say: E(t) = 0.625t + 68.95, just replace all Ys with E and Xs with t

OpenStudy (bahrom7893):

Now to find E(14): E(14) = 0.625*14 + 68.95 = 77.7

OpenStudy (bahrom7893):

so in 2007, the expectancy should be about 77.7 years

OpenStudy (anonymous):

thanks bahrom7893 so the linear function for E(t) = 0.625t + 68.95 is that correct

OpenStudy (bahrom7893):

yes

OpenStudy (anonymous):

ok than you u get a medal

OpenStudy (anonymous):

can i become your fan

OpenStudy (bahrom7893):

sure lol